OpenStudy (anonymous):

how do I find all zeroes of -3x^4+27x^2+1200=0

OpenStudy (anonymous):

divide by -3 would be a good start

OpenStudy (anonymous):

giving \(x^4-9x^2-400=0\) then you can try to factor, or else use the quadratic formula

OpenStudy (anonymous):

by some math teacher miracle this one actually factors as \[(x^2-25)(x^2+16)=0\] \[x^2=-16\] gives two complex zeros, \[x^2=25\] gives the two obvious real ones

OpenStudy (anonymous):

this is what I have -3(x-5)(x+5)(x^2+16)

OpenStudy (anonymous):

'math teacher miracle' indeed!

OpenStudy (anonymous):

do I set x-5=0 and x+5=0 but what about x^2+16=0 what would it be

OpenStudy (anonymous):

Yes. x^2 = -16 will give two imaginary solutions.

OpenStudy (anonymous):

=-i4

OpenStudy (anonymous):

You'll get a positive and negative 4i.